Hydrated Barium Chloride Trial One
Mass of empty crucible and cover 37.57 g
Mass of crucible, cover and hydrate 47.65 g
Mass after first heating 46.19 g
Mass after second heating 46.16 g
Mass of one mole of the anhydrous salt
2. For both trials, calculate the number moles of water removed by heat from your sample of hydrate.
1
Expert's answer
2020-05-08T14:10:28-0400

1. 47.65-46.19 = 1.46 g of first water.

2. 46.19-46.16 =0.03 g of second water.

BaCl2 137+71=208.